HW Help Save& Exit Submi Check my work 14-9 Issuance f bonds; effective interest amortization schedule; financial statement effects LO14-2] When Patey Pontoons issued 1 and maturity was 0% bonds on January 1, 2018, with a face amount of $880,000, the market yield for bonds of similar risk nterest is paid semian ly on June 11%. The bond s mature December 31 , 2021 (4 years). I nual 30 and December 31. of S1. PV of S1. FVA of S1, PVA of S1. EVAD of $1 and PVAD of $) (Use appropriate factorts) from the tables provided.) 1. Determine the price of the bonds at January 1, 2018. 2. Prepare the journal entry to record their issuance by Patey on January 1, 2018. an amortization schedule that determines interest at the effective rate each period. 4. Prepare the journal entry to record interest on June 30, 2018. 5. What is the amount related to the bonds that Patey will report in its balance sheet at December 31, 2018? 6. What is (Ignore income taxes) 7. Prepare the appropriate journal entries at maturity on December 31, 2021, the amount related to the bonds that Patey will report in its income statement for the year ended December 31, 2018? Complete this question by entering your answers in the tabs below. Req 2 Req 3 Req 4 Req S and 6 Req 7 e price of the bonds at January 1, 2018. (Round final answers to the nearest whole dollar.) < Prev 5 of 10 İİ Next >
Answer #1
1 Face amount $880,000
Pmt Semi annual interest $44,000 (880000*(0.1/2))
Rate Semi annual Market Yield =(11/2)%=5.5%               0.055
Nper Number of semi annual periods 8 (4*2)
Fv Maturity amount at end of 4 years $880,000
Price of Bond on January 1, 2018 $852,128 ( Using PV function of excel with Rate=0.055,Nper=8, Pmt=-44000, Fv=-880000)
2 JOURNAL ENTRY ON ISSUANCE
Date Account Title Debit Credit
.January1, 2018 Cash $852,128
Discount on Bond payable $27,872 (880000-852128)
Bond Payable $880,000
3 AMORTIZATION SCHEDULE
A B C D=C-B E F G=F-E
Date Interest Payment 5% *Face value Interest expense Amortization Debit Balance Credit Balance Book value
5.5%*Previous Book value in G of Bond Discount in Bond Discount in Bonds payable of the Bonds
.Jan1, 2018 - $27,872 $880,000 $852,128
.June30,2018 $44,000 $46,867 $2,867 $25,005 $880,000 $854,995
.Dec.31, 2018 $44,000 $47,025 $3,025 $21,980 $880,000 $858,020
.June30,2019 $44,000 $47,191 $3,191 $18,789 $880,000 $861,211
.Dec.31, 2019 $44,000 $47,367 $3,367 $15,423 $880,000 $864,577
.June30,2018 $44,000 $47,552 $3,552 $11,871 $880,000 $868,129
.Dec.31, 2020 $44,000 $47,747 $3,747 $8,124 $880,000 $871,876
.June30,2018 $44,000 $47,953 $3,953 $4,171 $880,000 $875,829
.Dec.31, 2021 $44,000 $48,171 $4,171 $0 $880,000 $880,000
Totals $352,000 $379,872 $27,872
4 JOURNAL ENTRY TO RECORD INTEREST ON JUNE 30,2018
Date Account Title Debit Credit
.June30,2018 Interest expense $46,867
Bond Discount $2,867
Cash $44,000